Pham can work as many or as few hours as she wants at the college bookstore for $11 per hour. But due to her hectic schedule, she has just 16 hours per week that she can spend working at either the bookstore or at other potential jobs. One potential job, at a café, will pay her $12 per hour for up to 6 hours per week. She has another job offer at a garage that will pay her $10 an hour for up to 5 hours per week. And she has a potential job at a daycare center that will pay her $8.50 per hour for as many hours as she can work. If her goal is to maximize the amount of money she can make each week, how many hours will she work at the bookstore?
1
Expert's answer
2018-05-22T15:43:08-0400
Pham can earn TR1 = $11*16 = $176 at the college bookstore.
At a café she can earn TR2 = $12*6 = $72.
At a garage she can earn TR3 = $10*5 = $50.
At a daycare center she can earn TR4 = $8.50*16 = $136.
If her goal is to maximize the amount of money she can make each week, then she should work 16 hours per week at the bookstore to earn $176.